OpenStudy (anonymous):

simplify the squareroot of (12x^10y^2)

OpenStudy (agreene):

\[\sqrt{12x^{10}y^{2}}=\sqrt{12}\sqrt{x^{10}y^2}=2\sqrt{3}\sqrt{x^{10}y^2}\]

OpenStudy (anonymous):

it can be further simplified as 2x^5y sq.root of (3x^5)

OpenStudy (agreene):

You can only break the radicals of x and y if they are positive.

OpenStudy (anonymous):

but if x and y are negative then their even powers make them p0ssitive isn't it?